# Folded Context Condensation in Path Integral Formalism for Infinite Context Transformers

### v4

@@ V-A Memory usage measurement & Speed Comparison
The Condensation model, represented by red lines in Figure 5, exhibits a linear increase in memory usage proportional to the sequence length. In contrast, the Llama model, shown with blue lines, deviates from this linear trend, indicating a more complex memory scaling behavior. At a total sequence length of 8K, the Llama model consumed 4.33 times more memory compared to our Condensation model with batch 1. Similarly, at 4K, the Llama model required 2.78 times more memory than the Condensation model with batch 4.
@@
The memory and speed experiments were conducted using the configurations shown in Table I, with both the Condensation and Llama models tested on a single NVIDIA H100 GPU. All memory experiments for the Condensation model were conducted using a segment length of 256. To compare the impact of the Condensation mechanism on memory usage and training speed, we used a reduced-scale model with 126 million trainable parameters.
@@ V-B Passkey retrieval task
The Passkey retrieval task embeds a passkey (a four-digit number) within a long document containing unrelated text and challenges the model to retrieve it. This task serves as a simple yet intuitive benchmark for evaluating whether the model retains critical information from previous segments or disregards it in favor of new input. In this study, we used the Llama-3.2-1B model as a baseline with Condensation logic applied, and the model was continuously pretrained on the Fine-web dataset [12] to adapt to the new logic with 512 segment length. The model was then finetuned on the Passkey task dataset starting with a sequence length of 2K, followed by incremental increases to 4K, 8K, and ultimately 16K, which was more efficient than directly finetuning to 16k. The model was finetuned up to 16k because of the constraints of memory resources. We evaluated the Condensation model further on 32K and 64K length datasets.
@@
The Fig. 8 depicts evaluation results for sequence lengths of 2K, 4K, 8K, 16K, and zero-shot evaluations at 32K and 64K. (Note that each results were evaluated with a 16K length fine-tuned model) The y-axis represents accuracy, while the x-axis indicates the depth, defined as the absolute position where the passkey is embedded. The model demonstrated the ability to effectively convey information across segments, achieving 100% accuracy within the sequence lengths on which it was fine-tuned. However, performance declined when evaluated on sequences longer than the model’s fine-tuning range. Enabling the model to retrieve passkeys beyond the finetuning range in a zero-shot manner is a potential avenue for future work.
@@ V-C Summarization task
While the passkey task is designed to evaluate the model’s ability to transfer information across segments, the summarization task assesses its capability to comprehend and integrate the full context. For this purpose, we scaled the model to Llama-3.2-3B and conducted continual pretraining on the Fine-web dataset [12] with 512 segment length. Subsequently, the model was first finetuned on the CNN/DailyMail dataset [16], which comprises a collection of news articles and their corresponding highlights. The highlights were treated as summaries, while the news articles served as the contextual input. To train the model on longer contexts, the Booksum dataset was filtered to include sequences of 5.5K length and combined with CNN/DailyMail for the second experiment. Two baseline models were prepared for benchmarking: Llama-3.2-3B-Instruct and Llama-3.2-3B fine-tuned exclusively on the CNN/DailyMail dataset. We employed a simple generation setting with a temperature of 0.3, generating 256 tokens without beam search.
@@ VI Conclusion
Our work introduces a quantum-inspired reformulation of Transformer architectures by leveraging Path Integral formalism to reinterpret the attention mechanism as a dynamic, temporal evolution of token states. By incorporating temporal dynamics into token state representations, the proposed Path Integral formulation provides a theoretical foundation for understanding transformer layers and attention mechanisms. This framework connects theoretical insights with practical architectural design, allowing constructive interference between tokens to participate in attention across segments. It also suggests that relatively compact token representations can effectively condense contextual information in a recurrent manner. Our experimental evaluations—including passkey retrieval and summarization tasks—demonstrate that this approach achieves both superior memory efficiency and competitive performance across extended sequence lengths. This perspective offers a promising direction for designing novel Transformer architectures and exploring quantum-inspired algorithms, opening up opportunities for future research. We plan to further expand on these ideas and share subsequent findings in future publications. We hope this work serves as a foundation for a deeper theoretical understanding of Transformer architectures and contributes to bridging concepts from quantum mechanics and artificial intelligence.
@@ Appendix A Passkey retrieval and Summarization task configuration
The table III presents the experimental configurations for pretraining and finetuning models on Passkey and Booksum tasks. For the Passkey task, a Condensation model based on Llama-3.2-1B with 16 layers and a 2048 embedding size was continuously pretrained on the Fineweb-edu dataset (filtered for sequences over 4K) and finetuned on the Passkey dataset, achieving a validation loss of 2.323 after continuous pretraining. For the Booksum task, a larger Condensation model based on Llama-3.2-3B with 28 layers and a 3076 embedding size was continuously pretrained on Fineweb-edu (filtered over 6K) and finetuned across two setups: (1) CNN/DailyMail, (2) CNN/DailyMail with filtered Booksum data. All experiments utilized the ZeRO-2 optimization [14], cosine learning rate decay, Bfloat16 mixed precision, and a global batch size of 64. We trained all models until we observed convergence of validation loss.
@@ Appendix B Summarization task results
Fig.  11 compares the performance of the Condensation 3B model fine-tuned on a 5.5K context length against the LLaMA-3.2-3B Instruction model. Due to memory constraints, we could not train an LLaMA model with a 5.5K context length for direct comparison. The results show that the Condensation model consistently outperforms the instruction-tuned baseline across all ROUGE metrics (ROUGE-1, ROUGE-2, ROUGE-L), demonstrating the effectiveness of the Condensation mechanism in handling long-document summarization tasks, even at extended context lengths.
Fig.  12 presents a comparison between two Condensation 3B models: one fine-tuned on CNN/DailyMail combined with filtered BookSum data using a 3K context length, and the other fine-tuned on the same dataset with additional data using a 5.5K context length. The results indicate that extending the context length to 5.5K during training does not lead to significant performance degradation when evaluated on the CNN/DailyMail dataset.

### v5

@@ V-A Memory usage measurement & Speed Comparison
The Condensation model, represented by red lines in Fig. 5, exhibits a linear increase in memory usage proportional to the sequence length. In contrast, the Llama model, shown with blue lines, deviates from this linear trend, indicating a more complex memory scaling behavior. At a total sequence length of 8 K (K represents 1024), the Llama model consumed 4.33 times more memory compared to our Condensation model with batch size of 1. Similarly, at 4 K, the Llama model required 2.78 times more memory than the Condensation model with batch size of 4.
@@
The memory and speed experiments were conducted using the configurations shown in Table I, with both the Condensation and Llama models tested on a single NVIDIA H100 GPU. In our experiments, we selected a segment size of 256 and 512 tokens for the Condensation model. This range was optimal given our resource constraints, requiring minimal additional GPU memory. While larger segment sizes could improve performance, we prioritized evaluating the model’s ability to transfer information across segments, necessitating multiple segment blocks. To compare the impact of the Condensation mechanism on memory usage and training speed, we used a reduced-scale model with 126 million trainable parameters.
@@ V-B Passkey retrieval task
The Passkey retrieval task embeds a passkey (a four-digit number) within a long document containing unrelated text and challenges the model to retrieve it. This task serves as a simple yet intuitive benchmark for evaluating whether the model retains critical information from previous segments or disregards it in favor of new input. In this study, we used the Llama-3.2-1B model (B represents Billion) as a baseline with Condensation logic applied, and the model was continuously pretrained on the Fine-web dataset [12] to adapt to the new logic with a 512-segment length. The model was then finetuned on the Passkey task dataset, starting with a sequence length of 2 K, followed by incremental increases to 4 K, 8 K, and ultimately 16 K, which was more efficient than directly finetuning to 16 K. The model was finetuned up to 16 K because of the constraints of memory resources. We evaluated the Condensation model further on 32 K and 64 K length datasets.
Fig. 8 depicts evaluation results for sequence lengths of 2 K, 4 K, 8 K, 16 K, and zero-shot evaluations at 32 K and 64 K. (Note that each results were evaluated with a 16 K length fine-tuned model) The y-axis represents accuracy, while the x-axis indicates the depth, defined as the absolute position where the passkey is embedded. The model demonstrated the ability to effectively convey information across segments, achieving 100% accuracy within the sequence lengths on which it was fine-tuned. However, performance declined when evaluated on sequences longer than the model’s fine-tuning range. Enabling the model to retrieve passkeys beyond the finetuning range in a zero-shot manner is a potential avenue for future work.
@@ V-C Summarization task
While the passkey task is designed to evaluate the model’s ability to transfer information across segments, the summarization task assesses its capability to comprehend and integrate the full context. For this purpose, we scaled the model to Llama-3.2-3B and conducted continual pretraining on the Fine-web dataset [12] with a 512-segment length. Subsequently, the model was first finetuned on the CNN/DailyMail dataset [16], which comprises a collection of news articles and their corresponding highlights. The highlights were treated as summaries, while the news articles served as the contextual input. To train the model on longer contexts, the Booksum dataset was filtered to include sequences of 5.5 K length and combined with CNN/DailyMail for the second experiment. Two baseline models were prepared for benchmarking: Llama-3.2-3B-Instruct and Llama-3.2-3B fine-tuned exclusively on the CNN/DailyMail dataset. We employed a simple generation setting with a temperature of 0.3, generating 256 tokens without beam search.
@@ VI Conclusion
[…]
By mapping the attention and neural structures of Transformer layers to the Path Integral framework, we identified a crucial challenge: maintaining phase coherence across token states to avoid destructive interference that leads to information loss. To address this, we introduced a periodic phase gap mechanism and recurrent segment processing, ensuring that token state evolution across segments preserves historical context through constructive interference. This mechanism is mathematically enforced by the temporal alignment condition, $t_{n}^{s}=t_{n}^{s-1}+\Delta t_{n}$, which makes an integer-multiple phase difference between segments at each layer, ensuring continuous information propagation and long-term dependency retention. This approach can be viewed as an artificial intelligence counterpart to human short-term memory, enhancing the quality of multi-turn conversations between humans and AI systems.
Experimentally, we validated our approach using two benchmark tasks—Passkey retrieval and long-document summarization. Our method demonstrated effective contextual retention across extended sequences, while notably scaling memory usage linearly rather than quadratically with sequence length, thus significantly enhancing computational efficiency compared with traditional Transformer models. Empirical evaluations confirmed that our model achieved comparable or superior performance in contextual information retrieval compared with the baseline Llama-3.2 models, while significantly reducing computational resources, as quantitatively evidenced by memory consumption and training speed metrics, where the conventional Transformer required up to 2.78 times more memory and 1.68 times more training time at 4 K sequence length (4.33 times more memory at 8 K).
In conclusion, our quantum-inspired formulation provides both a deeper theoretical understanding and substantial practical improvements in Transformer architectures. By leveraging Path Integral formalism, we not only reinforce interpretability but also substantially mitigate computational bottlenecks associated with long-range dependency modeling, opening promising avenues for future research into efficient and scalable Transformer-based models.
@@ Appendix A Passkey retrieval and Summarization task configuration
Table III presents the experimental configurations for pretraining and finetuning models on Passkey and Booksum tasks. For the Passkey task, a Condensation model based on Llama-3.2-1B with 16 layers and a 2048 embedding size was continuously pretrained on the Fineweb-edu dataset (filtered for sequences over 4 K) and finetuned on the Passkey dataset, achieving a validation loss of 2.323 after continuous pretraining. For the Booksum task, a larger Condensation model based on Llama-3.2-3B with 28 layers and a 3076 embedding size was continuously pretrained on Fineweb-edu (filtered over 6 K) and finetuned across two setups: (1) CNN/DailyMail, (2) CNN/DailyMail with filtered Booksum data. All experiments utilized the ZeRO-2 optimization [14], cosine learning rate decay, Bfloat16 mixed precision, and a global batch size of 64. We trained all models until we observed convergence of the validation loss.
@@ Appendix B Summarization task results
Fig. 11 compares the performance of the Condensation 3 B model fine-tuned on a 5.5 K context length against the Llama-3.2-3B Instruction model. Due to memory constraints, we could not train a Llama model with a 5.5 K context length for direct comparison. The results show that the Condensation model consistently outperforms the instruction-tuned baseline across all ROUGE metrics (ROUGE-1, ROUGE-2, ROUGE-L), demonstrating the effectiveness of the Condensation mechanism in handling long-document summarization tasks, even at extended context lengths.
Fig. 12 presents a comparison between two Condensation 3 B models: one fine-tuned on CNN/DailyMail combined with filtered BookSum data using a 3 K context length, and the other fine-tuned on the same dataset with additional data using a 5.5 K context length. The results indicate that extending the context length to 5.5 K during training does not lead to significant performance degradation when evaluated on the CNN/DailyMail dataset.
@@ Appendix D Summarization Task
Acknowledgment
This work was supported in part by the High Performance Computing program through the National IT Industry Promotion Agency (NIPA) and Korea Association for ICT Promotion (KAIT).